OpenStudy (anonymous):

Please help! Medal will be given. Simplify the expression. Rationalize all denominators. Assume that all variables are positive. Equation will be in another post.

OpenStudy (anonymous):

\[\sqrt[3]{-192a ^{7}b^3}\]

Nnesha (nnesha):

you can change root to exponent for example \[\huge\rm \sqrt[\color{red}{n}]{x^m} = x^\frac{ m }{ \color{red}{n} }\]

OpenStudy (anonymous):

@Nnesha - What would m be?

Nnesha (nnesha):

first factor -192

Nnesha (nnesha):

 one of number should be perfect cube root :)